Phyllis invested 35000 dollars, a portion earning a simple interest rate of 5 percent per year and the rest earning a simple interest rate of 6 percent per year. After one year the total interest earned on these investments was 1980 dollars. How much money did she invest at each rate?

Answer by stanbon(75887)

Phyllis invested 35000 dollars, a portion earning a simple interest rate of 5 percent per year and the rest earning a simple interest rate of 6 percent per year. After one year the total interest earned on these investments was 1980 dollars. How much money did she invest at each rate?
----------
Equation::
interest + interest = interest
0.05x + 0.06(35000-x) = 1980
5x + 6*35000 - 6x = 198000
----------------------------
-x = -12000
----
x = $12000 (amt. invested at 5%)
35000-x = $23000 (amt. invested at 6%)
-----------
Cheers,
Stan H.
=============